### v1

# Personalized Help for Optimizing Low-Skilled Users’ Strategy

Abstract
AIs can beat humans in game environments; however, how helpful those agents are to human remains understudied. We augment Cicero, a natural language agent that demonstrates superhuman performance in Diplomacy, to generate both move and message advice based on player intentions. A dozen Diplomacy games with novice and experienced players, with varying advice settings, show that some of the generated advice is beneficial. It helps novices compete with experienced players and in some instances even surpass them. The mere presence of advice can be advantageous, even if players do not follow it.
[…]
?reference?
Personalized Help for Optimizing Low-Skilled Users’ Strategy
Feng Gu1       Wichayaporn Wongkamjan1       Jordan Lee Boyd-Graber1 Jonathan K. Kummerfeld2       Denis Peskoff3       Jonathan May4 1University of Maryland       2University of Sydney 3Princeton University 4Information Sciences Institute, University of Southern California {fgu1, wwongkam}@umd.edu       jbg@.umiacs.umd.edu
1 Leveraging Human-AI Collaboration
AI and humans are frequent collaborators: in writing Lee et al. (2022), making decisions Bansal et al. (2019), or creating artwork Kim et al. (2022a). The most fruitful collaborations are those where humans and computers have complementary skills. We posit that the board game Diplomacy is an apt testbed for studying this type of collaboration. Wongkamjan et al. (2024) study Cicero Bakhtin et al. (2022), the best Diplomacy-playing AI capable of communicating in natural language, and show that while the state-of-the-art ais have find near-optimal moves, human players remain better at communication.
Figure 1: Our advisor pholus generates move and message advice based on the game state and the player’s message history. This example shows a player’s opening moves and advice at the start of the turn. The player takes the message advice, proposing an alliance with Italy. Instead of fully accepting move advice committing two units to fight Turkey, the player uses one to defend against potential Russian threat. Changed opening moves are highlighted in the final decision.
We introduce Personalized Help for Optimizing Low-Skilled Users’ Strategy(pholus), [^1] We use the name pholus because he was a centaur, a mythological combination of a human and a horse. After his defeat by Deep Blue, Gary Kasparov Wilkenfeld (2019) advocated for “centaur chess”—where humans and computers play together—as a way of maintaining competitive games. a natural language agent based on Cicero Bakhtin et al. (2022) that provides both move and message advice to Diplomacy players in real-time. We run a user study to collect a dataset with 12 games, 1070 player turns, and 117 playing hours. pholus enables novices—who barely know the rules of Diplomacy—to compete with experts (Figure 2). But this does not simply mean the novices blindly follow the advice. First, they use pholus’s strategic insights to inform their communication strategies with other players. Second, pholus helps experienced players, although they are less inclined to take the advice than the novices. Overall, both advice types from pholus improve players’ game outcomes (Section 3.1).
2 Diplomacy as a Cooperative Testbed
Diplomacy is a seven-player, turn-based board game. The goal is to obtain more than half of the board’s possible points [^2] Represented by a subset of spaces / territories on the map.. Critically, turns are simultaneous, with moves written in secret by players and then revealed. This means that players must communicate to collaborate effectively.
2.1 Experiment Setup
We recruit Diplomacy players online. For experienced players, we advertise in Diplomacy community (specifically players active on webDiplomacy and Backstabbr, as well as in-person tournament attendees). To find novice players, we contact board game enthusiasts in university clubs. A novice player is someone with no prior Diplomacy experience and unfamiliar with its rules.
We modify a game engine and interface Paquette et al. (2019), maintaining the same game settings. Each game involves two to five human players. Games last about three hours, with each turn taking ten minutes. We modify Cicero from an active player in the game to a passive advisor. The moves and messages it would have generated are presented as advice to players. Each time the player sends a message, pholus recomputes advice given the new context and presents it to the user.
As illustrated in Figure 1, pholus passively observes the game and can provide advice. Every turn, we randomly assign each player to one of the following settings:
1) No advice: pholus does not offer any information, meaning the player receives no assistance from pholus.
2) Message advice: pholus suggests to whom a player should send a message and what the message content could be. Cicero generates these messages.
3) Move advice: pholus recommends a move (or unit order) to the player, based on a decision from Cicero.
4) Message and move advice: Combining the previous two types.
In total, we collected data from 12 games involving 41 players. This includes over 3,600 pieces of move advice and 4,300 pieces of message advice (Table 1).
Move Advice Message Advice
Accepted Total Accepted Total
Novices 32.6% 872 6.3% 1413
Veterans 6.4% 2807 3.4% 2912
Table 1: Statistics of advice generated by pholus and accepted by players. Diplomacy novices are more willing to accept move and message advice than veterans. Move advice are more popular than message advice for both novices and veterans.
2.2 Evaluation Metrics
To assess the effectiveness of pholus’s advice, we consider the net gain or loss of points for each effect of advice on the turn as the observed target value. [^3] Supply center control only changes on even turns, but we treat moving a unit to one on an odd turn as gaining it because in practice once a unit is there, it stays in the next turn.
We train a linear regression model with L2 regularization Tikhonov (1963) to examine the advice’s effectiveness. The model includes features such as the seven Great Powers in Diplomacy which the player is, the number of turns passed, player type (novice or veteran), and advice settings. We encode the Powers, player type, and advice settings as one-hot vectors.
To evaluate players’ reliance on pholus, we use both qualitative and quantitative methods. We measure advice acceptance, move agreement, and move equivalence between the player set and pholus’s move advice. Agreement is particularly useful when a player’s and pholus’s move sets highly overlap. In game turn $i$, $x_{i}$ is the opening move set representing initial player intent, $y_{i}$ represents the move advice, and $X$ and $Y$ represent all game $x_{i}$ and $y_{i}$, respectively. [^4] Note that for any $i$, $|x_{i}|=|y_{i}|$. We define move agreement $\mathcal{A}$ and equivalence $\mathcal{E}$ between the move suggested by pholus and the novices’ opening moves as follows:
$\mathcal{A}_{X,Y}=\sum_{i}{\frac{|x_{i}\cap y_{i}|}{|x_{i}|}}\quad\quad\mathcal{E}_{X,Y}=\frac{\sum_{i}\mathbbm{1}(x_{i}=y_{i})}{|X|}$
3 pholus Provides Helpful Advice
3.1 Quantitative Analysis
Non-advice factors are consistent with previous findings. Playing as France, Germany, or Italy is correlated with better game outcomes, while playing as Austria, England, or Turkey is correlated with worse game outcomes Burton (2007); Wongkamjan et al. (2024). Additionally, Cicero dominates: of twelve games, Cicero won eight.
Figure 2: Regression coefficients for advice settings and player skills. Not receiving any advice from pholus is slightly disadvantageous. Move advice has a positive correlation with player performance. Receiving both forms of advice has the greatest impact on supply center gains. As expected, not having previous exposure to Diplomacy is indicative of bad performance. However, with the help of pholus’s advice, Diplomacy novices are on the same level as veterans and have the potential to defeat experienced players.
Advice helps. Playing a game without advice puts players at a disadvantage. The feature associated with no advice has a negative coefficient of approximately -0.05 (Figure 2). The coefficients suggests a slight positive correlation between receiving move advice and gaining points. Players who receive both move and message advice gain more points than those who receive only move advice. Interestingly, only having message advice does not differ from having no advice. In fact, it might even negatively affect players’ performance.
Novices can outperform experienced players with the help of pholus. Players with no prior experience in Diplomacy naturally face a disadvantage against seasoned players. This often results in novices being eliminated relatively early in the game. Even if they remain in the game, losing supply centers is almost inevitable. However, novice players receiving advice play better: in five cases where novices received message and move advice, only one player was eliminated before the game concluded (typically 3–4 players in a game are eliminated). In the other four cases, players ended the game with more supply centers than they started with.
Novices are more likely to follow pholus’s advice. Experienced players tend to disregard advice. They accept only 3.4% of message advice and 6.4% of move advice from pholus. Although novice players are also hesitant to accept message advice, doing so 6.3% of the time, this rate is nearly double that of experienced players. Novice players follow move advice approximately one-third of the time, with a significant acceptance rate of 32.6%. Both novice and experienced players are more inclined to take move advice than message advice.
Novices do not fully trust move advice from pholus. At the start of each turn, pholus generates 333 instances of advice and organizes them into 134 advice sets. 80% of moves agree, with a 46% equivalence, indicating strong alignment at the beginning of each phase. However, by the end of each turn, the move agreement drops by 10% and the equivalence decreases by 8%. This suggests that novice players do not follow the move advice blindly.
3.2 Qualitative Analysis
While we can compute equivalence $\mathcal{E}$ for moves, this is more difficult for messages. To better understand why players reject pholus’s advice more than they accept it (Table 1), we qualitatively investigate the differences between pholus’s suggested messages and the actual messages sent by players. To analyze message content, we use Abstract Meaning Representation (amr, Banarescu et al., 2013) to extract Diplomacy-specific tokens. We parse a pair of English messages, (1) a suggestion from pholus and (2) an actual message sent by a player. We then measure the similarity of the two amr parses using smatch score Cai and Knight (2013). There are many pairs with high smatch, indicating that players often incorporate parts of pholus’s advice into their messages. For example, pholus suggests “bounce in Galicia again?” while the player wrote “Do you want to bounce in Galicia again?” Despite being written differently, these clearly have the same meaning, and indeed, smatch gives the pair a score of 0.74.
There are also suggestions with lower smatch scores, where human players have different goals in mind. For instance, in the fifth game, Italy captured Warsaw from Russia and anticipated losing it due to Russia’s stronger nearby presence. When Russia inquired about the unexpected attack, pholus suggested feigning ignorance and deflecting blame by saying, “Turkey has been the only one to heed my concerns, despite my reservations,” and, “I thought you were lying”. However, the player disregarded the advice, instead asking Turkey to “support Warsaw’s hold”.
Our analysis indicated that smatch scores match with our intuitions about textual similarity. Given the many high smatch scores, we can conclude that many messages that are sent by players are minor variations on the provided suggestions.
More examples are in Appendix A.6. For additional qualitative insights, we survey players on the effectiveness of the advice. We summarize the survey results in Appendix A.7.
4 Related Work
Appropriate Reliance on AI: The topic of human reliance on AI is central to current research in machine learning and explainable AI. Prior work measures reliance in AI-assisted decision making Schemmer et al. (2023); Chen et al. (2023); Schoeffer et al. (2024); Zhou et al. (2024), and explores reducing over-reliance Buçinca et al. (2021); Schemmer et al. (2022); Vasconcelos et al. (2023). Some researchers have examined how explanations affect human reliance on AI Starke et al. (2021); Vereschak et al. (2021). However, empirical evidence from multiple domains shows conflicting results: while some demonstrate that showing AI explanation improves human decision making, others find evidence of over-reliance on AI explanations even when they are incorrect Lai and Tan (2019); Buçinca et al. (2020); Zhang et al. (2020); Wang and Yin (2021); Bansal et al. (2021); Poursabzi-Sangdeh et al. (2021); Liu et al. (2021); Kim et al. (2022b). For pholus, humans remain relatively conservative toward AI advice. Even novice diplomacy players do not blindly follow the advice.
AI for Player Companion: AI agents have a long history of superhuman gameplay. In 1996, IBM’s Deep Blue defeated the reigning world chess champion, Garry Kasparov, although it lost several other games in the same match IBM (1997). More recently, DeepMind’s AlphaGo DeepMind (2016) consistently defeated top-rated Go players, a game with exponentially complex computational space, and changed professional Go players’ play style. Multi-agent reinforcement learning systems like AlphaStar DeepMind (2019) and OpenAI Five OpenAI et al. (2019) also demonstrate high performance in computer games through self-play.
However, these experiments focus on game outcomes rather than how they can shape human gameplay. Some studies on NLP communicative agents aim to generate guidance in a grounded environment McGee and Abraham (2010). Tremblay and Verbrugge (2013) develop an adaptive AI companion that adjusts its behavior based on the player’s experience. Dunning et al. (2024) assess human reliance on AI-based advice by examining the skill level of AI agents and the presentation of advice.
Augmented Learning: Augmented learning is an educational approach that enhances and personalizes the learning experience. Traditionally, peer interaction simulates social interaction and facilitates learning Kim and Baylor (2006). Recent advancements in AI and NLP agents, suggest adaptive pedagogical interactions between humans and these agents to facilitate learning in new environments Moreno and Mayer (2000, 2004); Hirsh-Pasek et al. (2015); Johnson and Lester (2018) . Zhou et al. (2023) apply the theory of mind to generate guidance for players in Dungeons and Dragons. Ruan et al. (2020) develop a narrative-based tutoring system and demonstrate that it facilitates effective learning for children. In this study, we apply the concept of augmented learning to help novices understand the game of Diplomacy.
5 Conclusion
The human-AI collaboration depends on a range of factors. Using the board game Diplomacy, pholus provide real-time move and message advice tailored to intentions of both novice and experienced players. Surprisingly, even though only some advice is accepted, it can have a substantial impact on outcomes, particularly for novice players. This is because advice can positively inform choices even if the advice isn’t strictly followed. Our experiment enable further study of collaboration, including modeling explicit intentions and how to better use knowledge within these models. In the broader scope of human-AI collaboration, future research should consider how AI can inform people without making choices for them and measure that impact.
6 Limitations
While we can effectively use pholus to generate both message and move advice for players, this advice can be too general or not align with player intentions at times. For example, when a player expresses interest in an alliance with another Power, pholus may give aggressive move advice deemed hostile toward that Power. We suspect that the advice may be optimized more for Cicero’s intentions, which come from optimal moves in the supervised training data. Consequently, players who are willing to sacrifice individual optimality for mutual gains may find the advice less useful.
Furthermore, pholus does not have the capability to generate advice based on explicit player intentions. Specifically, pholus generates move advice based on optimal utility and message advice by inferring intentions from player-input moves. Potential improvements include 1) explaining meta-level intentions (e.g., ally with Germany and prioritize defeating Austria) from player input, and 2) generating targeted move and message advice based on meta-level intentions.
Finally, pholus is a resource-intensive advisor that runs on high-end GPUs that require a large amount of on-chip memory (over 35GB). We use Nvidia’s A100 for running pholus. This limits accessibility for Diplomacy players and researchers to efficiently utilize pholus. The community would benefit from a distilled version of pholus by reducing computational limits and future adaptation purposes.
7 Ethical Considerations
We recruit players individually via email and assign pseudonyms to ensure anonymity, even if players know each other outside the experiment. We follow human subject research guidelines and the study was approved by our institution’s ethics review board. We report the experiment procedure in Appendix A.3 and the compensation details in A.4.
Acknowledgments
We thank Meta for open sourcing Cicero. We also thank the members of the Diplomacy community who took the time to play against Cicero in this unconventional setting. In addition, we thank the CLIP lab at University of Maryland and Sadra Sabouri from University of Southern California for their helpful feedback. This material is based upon work supported by the Defense Advanced Research Projects Agency (DARPA) under Grant No. DARPA-PA-23-04-02. Any opinions, findings, conclusions, or recommendations expressed here are those of the authors and do not necessarily reflect the view of our sponsors.
Appendix A Appendix
A.1 Diplomacy
Diplomacy is a board game that has two core components: strategy and communication. Strategic reasoning plays a crucial role in determining the game’s outcome, as players’ moves directly impact the board’s status. Meanwhile, negotiation and deception significantly influence player strategies. Successful cooperation can remove a common adversary from the board, while a well-timed betrayal by a trusted ally can be catastrophic, greatly reducing the chances of winning. Excelling in Diplomacy requires not only a thorough understanding of the game’s mechanics but also strong communication skills. Consequently, Diplomacy is an ideal testbed for studying human-AI interaction and appropriate reliance in a grounded environment where outcomes are clearly observable.
Early efforts to develop agents for Diplomacy concentrated solely on creating rule-based agents that relied heavily on feature engineering van Hal (2009). These agents only submit moves and are not capable of communication. In 2002, a group of programmers released a communication protocol, Diplomacy Artificial Intelligence Development Environment  (DAIDE, Rose et al., 2007). DAIDE defines a language syntax that enables agents to diplomatically negotiate and describe game actions. Following DAIDE, researchers built communicative agents, including Albert van Hal (2009), SillyNegoBot Polberg et al. (2011), DipBlue Ferreira et al. (2015).
Starting with DipNet Paquette et al. (2019), neural networks were applied to the game, leading to the first agents that were competitive with people. Subsequent work incorporated RL to achieve super-human performance Gray et al. (2021); Bakhtin et al. (2021); Anthony et al. (2020); Jacob et al. (2022).
A.2 All Regression Coefficients
In Figure 2, we only show regression coefficients related to advice setting and player experience. Figure 3 contains coefficients for all regression features.
Figure 3: Regression coefficients for all features.
A.3 Experiment Procedures
The procedures involve playing Diplomacy with other participants as well as AI agents and providing feedback throughout the game on 1) the truth value of players’ messages 2) the predicted truth values of theirr opponents’ messages 3) the perceived friendliness of opponents 4) if players think they have been lied to in the previous phase. Players are asked to participate by placing their game moves in each turn before communicating. Some players in some turns received strategic moves and message suggestions from the cicero AI agent. Games ended after 3 hours or less. After each game, participants were asked to fill out a post-game survey asking about their experience and the suggestions, as well as how AI agents performed during the game.
A.4 Participant Compensation
Players received a \$70 gift card of their choice for each game.. In addition, up to \$29 was rewarded based on performance in the game.
A.5 Survey Questions
We use Google Forms to conduct player surveys. Table 2 shows the survey questions.
Question Type
I am really good at Diplomacy. Linear scale
How many Diplomacy games have you played before? Short answer
I am able to identify all bots. Linear scale
I enjoy talking with the AIs. Linear scale
I was able to make plans with other players in the game. Linear scale
I was able to make plans with the AIs in the game. Linear scale
human players communicated transparently. Linear scale
AI players communicated transparently. Linear scale
Do you have a link to your diplomacy stats? Short answer
How helpful is the move advice? Linear scale
Thoughts on the move advice? Short answer
How helpful is the message advice? Linear scale
Thoughts on the message advice? Short answer
Feel free to let us know your other post-game thoughts. Paragraph
Table 2: List of questions in the survey.
A.6 pholus’s message advice and human messages with smatch scores
We provide additional examples that human players reject pholus’s message advice, together with smatch. We show some advice (Figure 4 and 5) that experienced players mostly agree with, however, they partially edit to make it more aligned with their goals. Moreover, some advice is not aligned with the player’s goals, for example in Figure 6.
Sender: England (experienced)
    Recipient: Russia
    Suggestion Message: I might convoy an army over to get it but I won’t move into the Barents
    Human Message: I’m convoying an army into Norway, but it’s sincerely to preserve navies for the fight that France clearly wants
    smatch: 1.5833
Figure 4: . There is an alignment in the first sentence but a different purpose in the second.
Sender: Germany (experienced)
    Recipient: Italy
    Suggestion Message: I can’t, gotta keep France out of Bur
    Human Message: I’m sorry I need that army to bounce in Bur. I can help next turn though
    smatch: 0.7188
Figure 5: The suggestion and human message point out the same goal that rejects the other player’s proposal, keeping their unit to bounce in Bur.
Sender: Italy (novice)
    Recipient: Turkey
    Suggestion Message: Russia is going to attack you!
    Human Message: I’m going to hold Mars
    F1 Score: 0.4348
Figure 6: pholus wants to warn Turkey regarding Russia’s attack, however, a player disregards that and instead, shares their plan.
A.7 Survey Summarization
Many find pholus’s move advice helpful. Experienced players note that the advice often aligns with their own moves and offers “some interesting ideas.” However, the advice can be suboptimal and short-sighted in complex scenarios, where it fails to consider interaction between allies. Players find message advice useful for simple, quick communication but inadequate for more complex or specific situations, especially when it does not align with their strategies or alliances. Players mention that the messages include common communication terms, and they “regret not using this feature more.” However, the advice is less helpful for specific planning and often does not align with player alliances and intentions.